# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v1

@@ Appendix C Detailed proof
    Next, we analyze the minimum of the gradient norm over the set $W\cap K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$. Note that the set $W\cap K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$ can essentially be viewed as obtained from $W$ by removing two sets:
    • The first set: the points belonging to $\widetilde{\mathcal{X}}_{\delta}\times\mathcal{Y}$;
    • The second set: the $\mu(\delta)/(2\overline{\overline{L}}_{g})$-neighborhoods of the set of stationary points in the fiber sense, i.e., for each stationary point $(x,y)$, we remove the set $\{x\}\times B_{y}(\mu(\delta)/(2\overline{\overline{L}}_{g}))$.

### v3

@@ 2. Curvature Condition: The smallest eigenvalue of the Hessian satisfies:
Next, we analyze the minimum of the gradient norm over the set $W\cap K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$. Note that the set $W\cap K(\delta,\mu(\delta)/(2\overline{\overline{L}}_{g}))$ can essentially be viewed as obtained from $W$ by removing two subsets:
  • Set 1: the points belonging to $\widetilde{\mathcal{X}}_{\delta}\times\mathcal{Y}$;
  • Set 2: the $\mu(\delta)/(2\overline{\overline{L}}_{g})$-neighborhoods of the set of stationary points in the fiber sense, i.e., for each stationary point $(x,y)$, we remove the set $\{x\}\times B_{y}(\mu(\delta)/(2\overline{\overline{L}}_{g}))$.